A circular coffee table has a diameter of 5 ft. What will it cost to have the
top refinished if the company charges $3 per square foot for the refinishing

Answer by checkley77(12844)

A=PI*R^2
A=3.14(5/2)^2
A=3.15*2.5^2
A=3.14*6.25
A=19.625 FT^2 IS THE AREA.
3*19.624=$58.88 COST OF THE REFINISHING.